OpenStudy (anonymous):

Solve each equation or inequality: 18+7h^1\2=12

OpenStudy (anonymous):

it's just the h that is to the power of (1/2) right?

OpenStudy (anonymous):

7h^1/2 = -6 h^1/2 = -6/7 h = 36/49

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[18+7\sqrt{h}=12\]\[7\sqrt{h}=-6\]\[\sqrt{h}=-6/7\]\[h=36/49\]

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

\[\sqrt[3]{4m+1}-2=2\] what about this one?

OpenStudy (anonymous):

how would i do that?

OpenStudy (anonymous):

(4m +1)^(1/3) -2 = 2 (4m +1)^(1/3) = 4 ((4m+1)^(1/3))^3 = 4^3 4m + 1 = 64 4m = 63 m = 63/4

OpenStudy (anonymous):

thanks !